A criminologist developed a test to measure recidivism, where low scores indicated a lower probability of repeating the undesirable behavior. The test is normed so that it has a mean of 140 and a standard deviation of 36.

1.      What is the percentile rank of a score of 176 (4 pts)?

2.      What percentage of scores falls between 122 and 194 (4 pts)?

3.      Suppose an individual is in the 93.70th percentile in this test, what is his or her corresponding recidivism score (2 pts)?
Answer #1

a) P(x<176)

= P(z<(176-140)/36)

= P(z<1)

= 0.8413

b) P(122<x<194)

= P((122-140)/36<z<(194-140)/36)

= P(-0.5<z<1.5)

= 0.6247